OpenStudy (anonymous):

A metal blade spins at a constant rate of 5.0 revolutions per second about a pivot through one end of the blade. This rotation occurs in a region where the component of the earth's magnetic field perpendicular to the blade is 30 Micro T. If the blade is 60 cm in length, what is the magnitude of the potential difference between its ends?

OpenStudy (anonymous):

you can try to applie the fourth maxwell's equation or just try the lenz-Faraday's law ,you need to specify the blade's rotation axis in order to know the exact flux .(vertical or...)

OpenStudy (anonymous):

E =-d(NBAcos(theta)/dt? I don't know what to do..

OpenStudy (anonymous):

one revolution sweeps out an area of pi*.6^2 so dA/dt = 5pi*.6^2 emf = -B*dA/dt

OpenStudy (anonymous):

don't worry about the sign I guess, since it asks for magnitude..